OpenStudy (anonymous):

finding zeros : f(x)=49x^2-144 how will i solve this one?

OpenStudy (campbell_st):

what can you tell about the numbers 49 and 144...?

OpenStudy (anonymous):

perfect squares?

OpenStudy (campbell_st):

yep... so give you have 49x^2 this is the difference of 2 squares... thats why theres a subtraction sign this may help you factorise \[(ax +b)(ax - b)=a^2x^2 - b^2\] after factoring let it equal zero... then solve for x

OpenStudy (anonymous):

are there any shortcuts on how to solve it?

OpenStudy (campbell_st):

well you know 49 and 144 are square number.... so what do you think a and b might be...?

OpenStudy (anonymous):

7 and 12

OpenStudy (anonymous):

or is it 7x+12 7x-12 ??

OpenStudy (campbell_st):

thats it so you are solving (7x - 12)(7x + 12) = 0 so really you just need to solve 7x - 12 = 0 and 7x + 12 = 0 hope it makes sense

OpenStudy (anonymous):

so the roots are -7 and 12???

OpenStudy (campbell_st):

no, you need to solve 7x - 12 = 0 then its 7x = 12 x = 12/7 thats one solution. You need to do the other 7x + 12 = 0

OpenStudy (anonymous):

oohhhh

OpenStudy (campbell_st):

the solutions are the roots

OpenStudy (anonymous):

7x=-12 x= -12/7

OpenStudy (anonymous):

oh okay i understand it now thanks a lot... i have other problem here help me again please.. f(d)=2d^2-10d

OpenStudy (anonymous):

i don't know how to solve it because it's not a perfect square number

OpenStudy (skullpatrol):

Do you see a common factor?

OpenStudy (anonymous):

2?

OpenStudy (skullpatrol):

Yes, there is another one.

OpenStudy (campbell_st):

glad to help... lots of good people here to help with the next question

OpenStudy (anonymous):

2 and 2d?

OpenStudy (nirmalnema):

u can use a identity here a^2-b^2=(a+b)(a-b)

OpenStudy (skullpatrol):

$$\huge 2d^2-10d= 2d(d - 5)$$

OpenStudy (anonymous):

But where's the other squared number?

OpenStudy (anonymous):

I don't understand it :(

OpenStudy (nirmalnema):

49x^2-144 can be written as (7x+12)(7x-12) so x =12/7 and -12/7

OpenStudy (nirmalnema):

these are the two roots

OpenStudy (anonymous):

f(d)=2d^2-10d

OpenStudy (anonymous):

That question was answered already @nirmalnema

OpenStudy (anonymous):

Do you know how to do it now @MCath?

OpenStudy (nirmalnema):

ohh...ok

OpenStudy (anonymous):

i don't know this one f(d)=2d^2-10d

OpenStudy (skullpatrol):